Please show formula !!!Preparing a Cash Budget with Supporting Schedules using Excels Multi-Tab Cell Referencing and Basic Math Functions S&P Enter1 S & P Enterprises needs a cash budget for March. The following information is available. January February March 3 Data 4 Ac22 Budgeted selling & administrative depreciation in March $3,200 23 24 Other budgeted cash disbursements in March Equipment
Answer #1

Formulas are showing in the corresponding cell

A B C D
1 Expected Cash Collections : Amount $ Formula
2 March Cash Collections $          5,100 ='Budget Data'!D5
3 March Cash Collections on account
4 January Sales   ($25,000 * 22%) $          5,500 ='Budget Data'!B6*'Budget Data'!B12
5 February Sales ( $30,000 * 60%) $        18,000 ='Budget Data'!C6*'Budget Data'!B11
6 March Sales ( $40,000 * 15%) $          6,000 ='Budget Data'!D6*'Budget Data'!B10
7 Total Cash Collection $        34,600 SUM(B2:B6)
8
9 Payments to Suppliers Amount $
10 Accounts payable for inventory purchase ($10,500* 40%) $          4,200 ='Budget Data'!B15*'Budget Data'!B19
11 March Purchase ($23,500 * 60%) $        14,100 ='Budget Data'!B16*'Budget Data'!B18
12 Total Cash Payments $        18,300 SUM(B10:B11)
13
14 S & P Enterprises
15 Cash Budget
16 for month of March Amount $ Amount $ Formula
17
18 Cash Balance, March 1 $              11,500 ='Budget Data'!B29
19 Add: Cash Receipts
20 Collection from Customers $              34,600 =C7
21 Total Cash available before current financing $              46,100 =C18+C20
22 Less: Disbursements
23 Payment to suppliers $        18,300 =C12
24 Selling and administrative exps $        12,500 ='Budget Data'!B21
25 Equipment purchases $        14,000 ='Budget Data'!B25
26 Dividend Paid $          2,000 ='Budget Data'!B26
27 Total Disbursements $            (46,800) =SUM(B23:B26)
28 Excess (deficiency) of cash available for disbursements $                  (700) =C21+C27
29 Financing :
30 Borrowings $              10,700
31 Repayments $                       -  
32 Interest $                       -  
33 Total Financing $              10,700 =SUM(C30:C32)
34 Cash Balance, March 31 $              10,000 =C28+C33